The question is write an equation for the perpendicular bisector of the line segment joining the two points? the points are(-2, 2) and (5, 4)
Answer by mananth(16949)

mid point formula
(5-2)/2
3/2
(4+2)/2=3
(3/2,3)
x1 y1 x2 y2
-2 2 5 4

slope m = (y2-y1)/(x2-x1)
( 4 - 2 )/( 5 - -2 )
( 2 / 7 )
m= 2/ 7
The perpendicular will have slope of (-7/2) passes through (3/2,3)
m= -7/ 2

Plug value of the slope and point ( 3/ 2 , 3 ) in
Y = m x + b
3.00 = -21/4 + b
b= 3 - -21/4
b= 33/ 4
So the equation will be
Y = -7/2 x + 33/4
m.ananth@hotmail.ca